OpenStudy (anonymous):

The area of a rectangular painting is given by the trinomial x^2 + 4x – 21. What are the possible dimensions of the painting? Use factoring.

OpenStudy (anonymous):

A=x^2+7x-3x-21 =x(x+7)-3(x+7) =(x+7)(x-3) so L=(x+7) W=(x-3)

OpenStudy (hoblos):

x^2 + 4x – 21 = (x+7)(x-3) L=(x+7) W=(x-3)